OpenStudy (anonymous):

find an acute angle y such that sin y + cos y = sqrt 2

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

try \[\frac {\pi}{4}\]

OpenStudy (anonymous):

It works, but I don't really understand how to get to that conclusion on my own, besides just looking it up.

OpenStudy (anonymous):

i guess you are supposed to know that they are both \[\frac{\sqrt{2}}{2}\]there so when you add them up you get \[\sqrt{2}\]

OpenStudy (anonymous):

fair enough